# Age-Optimal Sampling and Routing under Intermittent Links and Energy Constraints

### v1

@@ Appendix B Proof of Lemma 5
Differentiating the action-value function given in (48) with respect to $y$, we obtain:
$\frac{\partial Q(y,r)}{\partial y}=\begin{cases}\lambda^{\star}-y,&\text{if }y<\lambda^{\star}-\mu_{r}\\ \mu_{r},&\text{if }y\geq\lambda^{\star}-\mu_{r}.\end{cases}$ (79)
For all $y$, the derivative is positive. Hence, $\forall r\in\mathcal{N}$, $Q(y,r)$ is monotonically increasing with $y$. As a result, for any $y_{2}\geq y_{1}$, we can establish that
$V^{*}(y_{2},\mathcal{R};\lambda^{\star})-V^{*}(y_{1},\mathcal{R};\lambda^{\star})$ (80)
$=\min_{r\in\mathcal{R}}Q(y_{2},r)-\min_{r\in\mathcal{R}}Q(y_{1},r)$
$\geq\min_{r\in\mathcal{R}}\{Q(y_{2},r)-Q(y_{1},r)\}\geq 0,$
which indicates that $V^{*}(y,\mathbf{l};\lambda^{\star})=\min_{r\in\mathcal{R}}Q(y,r)$ is monotonically increasing with $y$.
Since $\mu_{j}>\mu_{k}$, it follows that $\lambda^{\star}-\mu_{k}>\lambda^{\star}-\mu_{j}$. Then, using (79), we compute the difference:
$\frac{\partial Q(y,j)}{\partial y}-\frac{\partial Q(y,k)}{\partial y}$ (81)
[…]
In all cases, the difference is non-negative, thus
$\frac{\partial Q(y,j)}{\partial y}-\frac{\partial Q(y,k)}{\partial y}\geq 0,$ (82)
which completes the proof.
@@ Appendix C Proof of Lemma 6
Appendix D Proof of Lemma 7
As given by (79), the sole dependence of $\partial Q(y,i)/\partial y$ on $i$ is $\mu_{i}$. Thus, $\forall j,k\in\mathcal{G}_{\mu}$, we have
$\frac{\partial Q(y,j)}{\partial y}=\frac{\partial Q(y,k)}{\partial y},\forall y\in\mathbb{R}^{+},\forall j,k\in\mathcal{G}_{\mu}.$ (85)
Thus, if route $i$ is optimal at $y=0$, it is also optimal for every $y\in\mathbb{R}^{+}$:
$Q(y,i)=\min_{r\in\mathcal{G}_{\mu}}\{Q(y,r)\},\forall y\in\mathbb{R}^{+},$ (86)
which indicates that this route dominates the space $\mathcal{G}_{\mu}$. As a result, only one route from $\mathcal{G}_{\mu}$ will be included in $\mathcal{R^{\star}}$.
@@ Appendix F Proof of Lemma 3
Consider a policy $\pi=(k,w_{k},k,w_{k},\ldots)$ that selects a single route $k\in\mathcal{R}_{\infty}$ (always available) and waits for a constant time $w_{k}$ before generating the new sample after the previous update packet is delivered (i.e., $Z_{i}=w_{k},~\forall i$). We denote the long-term average age under this constant-wait route $k$ policy as $\lambda_{k}^{cw}$.
@@
Therefore, $\pi$ is a valid policy under the energy constraint. The long-term average age achieved by this policy can be calculated as
$\mathbb{E}_{Q_{k}}[Y]+\frac{\mathbb{E}_{Q_{k}}\left[(Y^{\prime}+w_{k})^{2}\right]}{2\mathbb{E}_{Q_{k}}[Y^{\prime}+w_{k}]}=\mu_{k}+\frac{\mu_{k}+w_{k}}{2}+\frac{\sigma_{k}^{2}}{2(\mu_{k}+w_{k})}.$ (89)
Let $\lambda^{\pi}$ denote the average age achieved by a policy $\pi\in\Pi$.
By definition of optimality, we have
$\lambda^{\star}\leq\lambda^{\pi},\quad\forall\pi\in\Pi.$ (90)
In particular, this implies:
$\lambda^{\star}\leq\min_{k\in\mathcal{R}_{\text{s}}}\lambda_{k}^{cw}=\min_{k\in\mathcal{R}_{\text{s}}}\frac{3\mu_{k}+w_{k}}{2}+\frac{\sigma_{k}^{2}}{2(\mu_{k}+w_{k})}.$ (91)
Since the AoI $\Delta(t)$ is non-negative, $\lambda^{\star}\geq 0$. This completes the proof.

### v2

@@ Appendix B Proof of Lemma 5
We start with the general form of the action-value function $Q(y,r)$. In the non-linear setting, the accumulated age penalty is captured by the primitive function $F(t)\triangleq\int_{0}^{t}f(\tau)d\tau$. The simplified action-value function, substituting the optimal waiting policy $z^{\star}=(\beta^{\star}(r)-y)^{+}$, is given by:
$Q(y,r)=$ $C_{r}-(\lambda^{\star}+cE_{\max})(\beta^{\star}(r)-y)^{+}$ (59)
$+\mathbb{E}_{Y^{\prime}\sim Q_{r}}\left[F(y+(\beta^{\star}(r)-y)^{+}+Y^{\prime})-F(y)\right],$
where $C_{r}$ collects terms independent of $y$ (such as the value function expectation and energy costs). We analyze the partial derivative $\frac{\partial Q(y,r)}{\partial y}$ in two regions defined by the threshold $\beta^{\star}(r)$.
Case 1: Waiting Region ($y<\beta^{\star}(r)$). In this region, the waiting time is positive: $z^{\star}=\beta^{\star}(r)-y$. Substituting this into (59), the linear term becomes $(\lambda^{\star}+cE_{\max})(\beta^{\star}(r)-y)$. The argument of the penalty function becomes $y+(\beta^{\star}(r)-y)+Y^{\prime}=\beta^{\star}(r)+Y^{\prime}$, which is constant with respect to $y$. Thus, the equation simplifies to:
$Q(y,r)$ $=C_{r}-(\lambda^{\star}+cE_{\max})(\beta^{\star}(r)-y)$ (60)
$+\mathbb{E}_{Y^{\prime}}[F(\beta^{\star}(r)+Y^{\prime})-F(y)].$
Differentiating with respect to $y$:
$\frac{\partial Q(y,r)}{\partial y}=(\lambda^{\star}+cE_{\max})-f(y).$ (61)
Recall the definition of the threshold $\beta^{\star}(r)=\inf\{t\geq 0:\mathbb{E}[f(t+Y^{\prime})]\geq\lambda^{\star}+cE_{\max}\}$. Since $\beta^{\star}(r), yy$ is strictly smaller than the infimum. This implies that $y$ does not satisfy the condition of the set. Therefore, the expected penalty at $y$ must be strictly less than the system cost:
$\mathbb{E}[f(y+Y^{\prime})]<\lambda^{\star}+cE_{\max}.$ (62)
Furthermore, since $Y^{\prime}\geq 0$ and $f(\cdot)$ is monotonically non-decreasing, we have $f(y)\leq\mathbb{E}[f(y+Y^{\prime})]$. Combining these inequalities yields:
$f(y)\leq\mathbb{E}[f(y+Y^{\prime})]<\lambda^{\star}+cE_{\max}.$ (63)
Consequently, $(\lambda^{\star}+cE_{\max})-f(y)>0$, proving the derivative is strictly positive in this region.
Case 2: Sampling Region ($y\geq\beta^{\star}(r)$). In this region, the waiting time is zero: $z^{\star}=0$. The linear term vanishes, and the equation simplifies to:
$Q(y,r)=C_{r}+\mathbb{E}_{Y^{\prime}}[F(y+Y^{\prime})-F(y)].$ (64)
Differentiating with respect to $y$:
$\frac{\partial Q(y,r)}{\partial y}=\mathbb{E}_{Y^{\prime}}[f(y+Y^{\prime})]-f(y).$ (65)
Since $Y^{\prime}\geq 0$ and $f(\cdot)$ is a monotonic non-decreasing function, it holds that $\mathbb{E}_{Y^{\prime}}[f(y+Y^{\prime})]\geq f(y)$. Thus, the derivative is always non-negative.
Conclusion: Combining both cases, we have $\frac{\partial Q(y,r)}{\partial y}\geq 0$ for all $y\in\mathbb{R}^{+}$. This proves statement (1). Since $V^{*}(y,\mathbf{l})=\min_{r\in\mathcal{R}(\mathbf{l})}Q(y,r)$ is the minimum of monotonically increasing functions, it is also monotonically increasing with $y$, proving statement (2).
@@ Appendix C Proof of Lemma 6
The target threshold is defined as $\beta^{\star}(r)=\inf\{t\geq 0:\Psi_{r}(t)\geq\lambda^{\star}+cE_{\max}\}$.
First, we determine the ordering of the thresholds. Consider any $t$ that satisfies the threshold condition for route $k$, i.e., $\Psi_{k}(t)\geq\lambda^{\star}+cE_{\max}$. By the ordering condition (15), we have $\Psi_{j}(t)\geq\Psi_{k}(t)$ for all $t$. Therefore, inequality $\Psi_{j}(t)\geq\lambda^{\star}+cE_{\max}$ holds. This implies that any time $t$ feasible for the infimum of route $k$ is also feasible for route $j$. Since the set of feasible times for route $k$ is a subset of that for route $j$, the infimum of the larger set must be smaller or equal:
$\beta^{\star}(j)\leq\beta^{\star}(k).$ (66)
Next, we evaluate the partial derivative of the action-value function with respect to $y$. From Appendix B, the derivative is given by:
$\frac{\partial Q(y,r)}{\partial y}=\begin{cases}\lambda^{\star}+cE_{\max}-f(y),&y<\beta^{\star}(r)\\ \Psi_{r}(y)-f(y),&y\geq\beta^{\star}(r)\end{cases}$ (67)
We analyze the difference $\Delta_{jk}(y)\triangleq\frac{\partial Q(y,j)}{\partial y}-\frac{\partial Q(y,k)}{\partial y}$ across the three intervals defined by $\beta^{\star}(j)\leq\beta^{\star}(k)$:
  • Region 1 ($y<\beta^{\star}(j)$): Both routes are in the waiting region.
    $\Delta_{jk}(y)=(\lambda^{\star}+cE_{\max}-f(y))-(\lambda^{\star}+cE_{\max}-f(y))=0.$ (68)
  • Region 2 ($\beta^{\star}(j)\leq y<\beta^{\star}(k)$): Route $j$ is in the sampling region, while route $k$ is in the waiting region.
    $\Delta_{jk}(y)$ $=(\Psi_{j}(y)-f(y))-(\lambda^{\star}+cE_{\max}-f(y))$ (69)
    $=\Psi_{j}(y)-(\lambda^{\star}+cE_{\max}).$
    Since $y\geq\beta^{\star}(j)$ and $\Psi_{j}(y)$ is non-decreasing, by the definition of the infimum, we must have $\Psi_{j}(y)\geq\lambda^{\star}+cE_{\max}$. Thus, $\Delta_{jk}(y)\geq 0$.
  • Region 3 ($y\geq\beta^{\star}(k)$): Both routes are in the sampling region.
    $\Delta_{jk}(y)$ $=(\Psi_{j}(y)-f(y))-(\Psi_{k}(y)-f(y))$ (70)
    $=\Psi_{j}(y)-\Psi_{k}(y).$
    By the assumed ordering condition (15), $\Psi_{j}(y)\geq\Psi_{k}(y)$ holds for all $y$, so $\Delta_{jk}(y)\geq 0$.
Combining these cases, we conclude that $\frac{\partial Q(y,j)}{\partial y}\geq\frac{\partial Q(y,k)}{\partial y}$ for all $y\in\mathbb{R}^{+}$.
Appendix D Proof of Lemma 7
@@ Appendix F Proof of Lemma 3
Consider a stationary policy $\pi=(k,w_{k},k,w_{k},\ldots)$ that consistently selects a single route $k\in\mathcal{R}_{\infty}$ (where $\mathcal{R}_{\infty}$ denotes the set of routes that are always available) and waits for a constant time $w_{k}$ before generating a new sample after the previous update is delivered.
@@
Thus, $\pi$ is a valid policy.
Next, we derive the long-term average age penalty $\lambda_{k}^{cw}$ achieved by this policy using the Renewal Reward Theorem. A renewal cycle $i$ begins immediately after the delivery of packet i1i-1. The age is reset to $Y_{i-1}$ (the delay of the previous packet). The system waits for $w_{k}$, and then transmits packet $i$, which experiences delay $Y_{i}$. The cycle ends when packet $i$ is delivered at time $t=w_{k}+Y_{i}$.
The accumulated penalty during one cycle is the integral of the penalty function $f(\Delta(t))$. The instantaneous age starts at $\Delta(0)=Y_{i-1}$ and grows linearly to $Y_{i-1}+w_{k}+Y_{i}$ at the moment of delivery. Thus, the total penalty area for cycle $i$ is:
$R_{i}=\int_{0}^{w_{k}+Y_{i}}f(Y_{i-1}+t)\,dt.$ (75)
Using the substitution $u=Y_{i-1}+t$ and the primitive function $F(t)\triangleq\int_{0}^{t}f(\tau)d\tau$, this integral becomes:
$R_{i}=F(Y_{i-1}+w_{k}+Y_{i})-F(Y_{i-1}).$ (76)
Taking the expectation over the i.i.d. delays $Y_{i-1},Y_{i}\sim Q_{k}$, the average penalty is:
$\lambda_{k}^{cw}=\frac{\mathbb{E}_{Y,Y^{\prime}\sim Q_{k}}[F(Y+w_{k}+Y^{\prime})-F(Y)]}{\mu_{k}+w_{k}}.$ (77)
By the definition of optimality, $\lambda^{\star}\leq\lambda^{\pi}$ for any valid policy $\pi$. Therefore:
$\lambda^{\star}\leq\min_{k\in\mathcal{R}_{\infty}}\lambda_{k}^{cw}$ (78)
$=\min_{k\in\mathcal{R}_{\infty}}\left(\frac{\mathbb{E}_{Y,Y^{\prime}\sim Q_{k}}[F(Y^{\prime}+w_{k}+Y)-F(Y^{\prime})]}{\mu_{k}+w_{k}}\right).$
Finally, for the lower bound, since the age $\Delta(t)\geq 0$ and $f(\cdot)$ is monotonic, the instantaneous penalty is lower-bounded by $f(0)$. Consequently, the long-term average penalty satisfies $\lambda^{\star}\geq f(0)$. This completes the proof.